Marvel Parts, Inc., manufactures auto accessories. One of the company’s products is a set of seat covers that can be adjusted to fit nearly any small car. The company has a standard cost system in use for all of its products. According to the standards that have been set for the seat covers, the factory should work 1,060 hours each month to produce 2,120 sets of covers. The standard costs associated with this level of production are: Total Per Set of Covers Direct materials $ 43,460 $ 20.50 Direct labor $ 9,540 4.50 Variable manufacturing overhead (based on direct labor-hours) $ 4,664 2.20 $ 27.20 During August, the factory worked only 500 direct labor-hours and produced 2,200 sets of covers. The following actual costs were recorded during the month: Total Per Set of Covers Direct materials (8,000 yards) $ 44,000 $ 20.00 Direct labor $ 10,340 4.70 Variable manufacturing overhead $ 5,500 2.50 $ 27.20 At standard, each set of covers should require 2.50 yards of material. All of the materials purchased during the month were used in production.

1. Compute the materials price and quantity variances for August. (Input all amounts as positive values. Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e, zero variance).)



2. Compute the labor rate and efficiency variances for August. (Input all amounts as positive values. Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e, zero variance).)


3. Compute the variable overhead rate and efficiency variances for August. (Input all amounts as positive values. Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e, zero variance).)
Answer #1

1) Material price variance = (8.2*8000-44000) = 21600 F

Material quantity variance = (2200*2.5-8000)*8.2 = 20500 U

2) Labor rate variance = (9*500-10340) = 5840 U

Labor efficiency variance = (1100-500)*9 = 5400 F

3) Variable overhead rate variance = (4.4*500-5500) = 3300 U

Variable overhead efficiency variance = (1100-500)*4.4 = 2640 F